# Lusztig sheaves, characteristic cycles and the Borel-Moore homology of Nakajima’s quiver varieties

### v1

@@ 5.1.2. Nakajima’s quiver variety
An element $(x,y,z)\in\mathbf{E}_{\mathbf{V}\oplus\mathbf{W},H^{(1)}}$ is said to be stable, if the zero space is the unique $I$-graded subspace $\mathbf{V}^{\prime}$ of $\mathbf{V}$ such that $x_{h}(\mathbf{V}^{\prime}_{h^{\prime}})\subset\mathbf{V}^{\prime}_{h^{\prime\prime}}$ and $y_{i}(\mathbf{V}^{\prime}_{i})=0$ for any $h\in H,i\in I$, see [22, Lemma 3.8]. We define $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)^{s}$ to be the subset of $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)$ consisting of stable elements, and define $\Lambda_{\mathbf{V},\mathbf{W}}^{s}$ to be the subset of $\Lambda_{\mathbf{V},\mathbf{W}}$ consisting of stable elements.
Notice that $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)^{s}$ and $\Lambda_{\mathbf{V},\mathbf{W}}^{s}$ are $\mathbf{G}_{\mathbf{V}}$-invariant. By [22, Lemma 3.10], the group $\mathbf{G}_{\mathbf{V}}$ acts freely on $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)^{s}$ and $\Lambda_{\mathbf{V},\mathbf{W}}^{s}$. Nakajima’s quiver varieties
@@ 5.1.3. Tensor product variety
acts on $\mathfrak{M}(\nu,\omega)$. By [24, Lemma 3.2], the $\lambda(\mathbb{G}_{m})$-fixed point set $\mathfrak{M}(\nu,\omega)^{\lambda(\mathbb{G}_{m})}$ is isomorphic to
@@
$\xi x^{\prime}=x,y\xi=y^{\prime},\xi z^{\prime}=z.$
@@
where $\mathrm{sw}:\mathfrak{M}(\nu^{\prime},\omega)\times\mathfrak{M}(\nu,\omega)\rightarrow\mathfrak{M}(\nu,\omega)\times\mathfrak{M}(\nu^{\prime},\omega)$ the natural map exchanging the coordinates. We remark that the sign twists in [22] is incorrect and one should use the sign twists in formulas of [23, Section 9.3.2]. The fundamental class $[\Delta(\nu,\omega)]$ of the diagonal defines operators
@@
By [24, Proposition 5.5], since ~$\tilde{\mathfrak{Z}}_{1}$ is a vector bundle over $\coprod\limits_{\nu}\mathfrak{L}(0,\omega^{1})\times\mathfrak{L}(\nu,\omega^{2})$, there is an $\mathbf{U}_{1}^{+}(\mathfrak{g})$-linear isomorphism given by Thom isomorphism $\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\tilde{\mathfrak{Z}}_{1},\mathbb{Q})\cong\mathbf{H}(\mathfrak{L}(\omega^{2})).$ Moreover, the subspace $\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\tilde{\mathfrak{Z}}_{1},\mathbb{Q})$ generates the $\mathbf{H}(\tilde{\mathfrak{Z}}(\omega))$ as a $\mathfrak{g}$-module. Nakajima raised following conjecture and proved that this conjecture holds for ADE case, see [24, Theorem 5.9].
Conjecture 5.3 ([24, Conjecture 5.6]).
@@ 6.1.1. Singular supports of $\mathcal{P}_{\nu}$
We refer [14, Section 12 and 13] for details of this subsection.
Proposition 6.1 ([14, Corollary 13.6]).
@@
such that $\epsilon_{i}(L)=\epsilon_{i}(\Phi(L)),\epsilon^{\ast}_{i}(L)=\epsilon^{\ast}_{i}(\Phi(L))$ for any $i\in I$, see subsection 3.6 for the definitions of $\epsilon_{i}(L),\epsilon^{\ast}_{i}(L)$. Moreover, by [12, Theorem 6.6.6], we have
@@
where the union runs over all $Z\in\Lambda_{\mathbf{V}}$ such that $\epsilon_{i}(Z)\geqslant\epsilon_{i}(L),\epsilon^{\ast}_{i}(Z)\geqslant\epsilon^{\ast}_{i}(L)$ for any $i\in I$. By [12, Lemma 8.2.1] we have $\rho_{i,t}(\Phi(L))=\Phi(L^{\prime})$ if and only if $\pi_{i,t}(L)=L^{\prime}$, and $\rho^{\ast}_{i,t}(\Phi(L))=\Phi(L^{\prime})$ if and only if $\pi^{\ast}_{i,t}(L)=L^{\prime}.$
@@ 6.1.3. Singular supports of $\mathcal{P}_{\nu,\omega^{1},\omega^{2}}$
In order to show $SS(L)$ is contained in $\Pi_{\mathbf{V},\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$, it suffices to check for $L=L_{\boldsymbol{\nu}^{\prime}\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}}$. By applying the proof of [14, Theorem 13.3 and Corollary 13.6] to $Q^{(2)}$, we can see that any point in $SS(L_{\boldsymbol{\nu}^{\prime}\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}})$ admits a flag $f$ of flag type $\boldsymbol{\nu}^{\prime}\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}$. In the flag $f$, there are two special subspaces. The first one is $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{2}$ corresponding to $\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}$, the second one is $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}$ corresponding to $\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}$. In particular, $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}$ is $(x,y^{1},z^{1},y^{2},z^{2})$-stable, and so $(z^{1}+z^{2})(\mathbf{W})\subset\mathbf{V}^{\prime\prime}$. Similarly, $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{2}$ is $(x,y^{1},z^{1},y^{2},z^{2})$-stable, and so $(y^{1}+y^{2})(\mathbf{V}^{\prime\prime})\subset\mathbf{W}^{2}$. Notice that $z^{2}=0$ implies that $(z^{1}+x^{2})(\mathbf{W}^{2})=0$. We get the proof. ∎
@@ 6.3. Characteristic cycles for framed quivers
By Proposition 6.4, $\mathcal{Q}_{\nu,\omega^{1}}$ is a full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}})$, hence $\mathcal{K}(\omega^{1})_{-1}^{\psi}$ can be regarded as a sub $\mathbb{Z}$-module of $\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}}))$. Compose the characteristic cycle map with $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}]$, we obtain a morphism
@@
here $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}]$ is the linear map $\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}}))\rightarrow\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}}))$ induced by the functor $\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}:L\mapsto L[{\rm{dim}}\mathbf{G}_{\Omega}]$.
@@
(7) $\mathbf{Ind}^{(r)}_{i}\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}\cong\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}\mathcal{F}^{(r)}_{i},$
and $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbf{G}_{\Omega}}]$ intertwines the operator $[\mathcal{F}^{(r)}_{i}]$ and the operator $[\mathbf{Ind}^{(r)}_{i}]$. Here we denote $\mathbf{Ind}$ by $\mathbf{Ind}^{(r)}_{i}$, in order to emphasize the change of dimension vectors.
Let $\bar{F}_{\mathbf{W}^{1}}$ be the subset of $T^{\ast}\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}$ consisting of $(x,y,z)$ such that $(x,y,z)(\mathbf{V}^{\prime\prime})\subseteq\mathbf{V}^{\prime\prime}$ and denote $\bar{F}_{\mathbf{W}^{1}}\cap\Lambda_{\mathbf{V},\mathbf{W}^{1}}$ by $\bar{F}^{nil}_{\mathbf{W}^{1}}$. By [9, Lemma 9.4], we have $\mathbf{G}_{\mathbf{V}}\times^{P}\bar{F}_{\mathbf{W}^{1}}\cong T^{\ast}_{W}(X\times X^{\prime})$, and we also have the following commutative diagram,
@@
hence the induction operator $\mathbf{Ind}^{(r)}_{i}:\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\Lambda_{\mathbf{V}^{\prime\prime},\mathbf{W}^{1}},\mathbb{Z})\rightarrow\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\Lambda_{\mathbf{V},\mathbf{W}^{1}},\mathbb{Z})$ is well-defined. By [9, Theorem 9.11], the top degree homology groups of CoHA considered in [28] and [5] realize the nilpotent enveloping algebra. Hence together with $[\mathbf{Ind}^{(r)}_{i}]^{\psi}$, the homology group $\bigoplus\limits_{\nu\in\mathbb{N}I}\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\Lambda_{\mathbf{V},\mathbf{W}^{1}},\mathbb{Z})$ becomes a ${}_{\mathbb{Z}}\mathbf{U}^{-}(\mathfrak{g})$-module. We have the following lemma.
@@
is a $\mathbf{U}^{-}(\mathfrak{g})$-linear map.
@@
By equation (7) and Theorem 2.2, the map ${\rm{CC}}^{\omega^{1}}$ intertwines the operator $[\mathcal{F}^{(r)}_{i}]$ and the operator $[\mathbf{Ind}^{(r)}_{i}]$. Since we use the same $\psi$-twist, the proposition is proved. ∎
@@
by using functorial properties of characteristic cycles, where $\mathfrak{P}^{(r)}_{i}$ is the generalization of the Hecke correspondence in [23, Section 5.3]. Then ${\rm{CC}}^{s,\omega^{1}}$ is ${}_{\mathbb{Z}}\mathbf{U}(\mathfrak{g})$-linear directly, but we do not need such strong properties here.

### v3

@@ 5.1.2. Nakajima’s quiver variety
An element $(x,y,z)\in\mathbf{E}_{\mathbf{V}\oplus\mathbf{W},H^{(1)}}$ is said to be stable, if the zero space is the unique $I$-graded subspace $\mathbf{V}^{\prime}$ of $\mathbf{V}$ such that $x_{h}(\mathbf{V}^{\prime}_{h^{\prime}})\subset\mathbf{V}^{\prime}_{h^{\prime\prime}}$ and $y_{i}(\mathbf{V}^{\prime}_{i})=0$ for any $h\in H,i\in I$, see [26, Lemma 3.8]. We define $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)^{s}$ to be the subset of $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)$ consisting of stable elements, and define $\Lambda_{\mathbf{V},\mathbf{W}}^{s}$ to be the subset of $\Lambda_{\mathbf{V},\mathbf{W}}$ consisting of stable elements.
Notice that $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)^{s}$ and $\Lambda_{\mathbf{V},\mathbf{W}}^{s}$ are $\mathbf{G}_{\mathbf{V}}$-invariant. By [26, Lemma 3.10], the group $\mathbf{G}_{\mathbf{V}}$ acts freely on $\mu_{\mathbf{V}\oplus\mathbf{W}}^{-1}(0)^{s}$ and $\Lambda_{\mathbf{V},\mathbf{W}}^{s}$. Nakajima’s quiver varieties
@@ 5.1.3. Tensor product variety
acts on $\mathfrak{M}(\nu,\omega)$. By [28, Lemma 3.2], the $\lambda(\mathbb{G}_{m})$-fixed point set $\mathfrak{M}(\nu,\omega)^{\lambda(\mathbb{G}_{m})}$ is isomorphic to
@@
$\xi x^{\prime}=x\xi,y\xi=y^{\prime},\xi z^{\prime}=z.$
@@
where $\mathrm{sw}:\mathfrak{M}(\nu^{\prime},\omega)\times\mathfrak{M}(\nu,\omega)\rightarrow\mathfrak{M}(\nu,\omega)\times\mathfrak{M}(\nu^{\prime},\omega)$ the natural map exchanging the coordinates. We remark that the sign twists in [26] is incorrect and one should use the sign twists in formulas of [27, Section 9.3.2]. The fundamental class $[\Delta(\nu,\omega)]$ of the diagonal defines operators
@@
By [28, Proposition 5.5], since ~$\tilde{\mathfrak{Z}}_{1}$ is a vector bundle over $\coprod\limits_{\nu}\mathfrak{L}(0,\omega^{1})\times\mathfrak{L}(\nu,\omega^{2})$, there is an $\mathbf{U}_{1}^{+}(\mathfrak{g})$-linear isomorphism given by Thom isomorphism $\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\tilde{\mathfrak{Z}}_{1},\mathbb{Q})\cong\mathbf{H}(\mathfrak{L}(\omega^{2})).$ Moreover, the subspace $\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\tilde{\mathfrak{Z}}_{1},\mathbb{Q})$ generates the $\mathbf{H}(\tilde{\mathfrak{Z}}(\omega))$ as a $\mathfrak{g}$-module. Nakajima raised following conjecture and proved that this conjecture holds for ADE case, see [28, Theorem 5.9].
Conjecture 5.3 ([28, Conjecture 5.6]).
@@ 6.1.1. Singular supports of $\mathcal{P}_{\nu}$
We refer [18, Section 12 and 13] for details of this subsection.
Proposition 6.1 ([18, Corollary 13.6]).
@@
such that $\epsilon_{i}(L)=\epsilon_{i}(\Phi(L)),\epsilon^{\ast}_{i}(L)=\epsilon^{\ast}_{i}(\Phi(L))$ for any $i\in I$, see subsection 3.6 for the definitions of $\epsilon_{i}(L),\epsilon^{\ast}_{i}(L)$. Moreover, by [16, Theorem 6.6.6], we have
@@
where the union runs over all $Z\in\Lambda_{\mathbf{V}}$ such that $\epsilon_{i}(Z)\geqslant\epsilon_{i}(L),\epsilon^{\ast}_{i}(Z)\geqslant\epsilon^{\ast}_{i}(L)$ for any $i\in I$. By [16, Lemma 8.2.1] we have $\rho_{i,t}(\Phi(L))=\Phi(L^{\prime})$ if and only if $\pi_{i,t}(L)=L^{\prime}$, and $\rho^{\ast}_{i,t}(\Phi(L))=\Phi(L^{\prime})$ if and only if $\pi^{\ast}_{i,t}(L)=L^{\prime}.$
@@ 6.1.3. Singular supports of $\mathcal{P}_{\nu,\omega^{1},\omega^{2}}$
In order to show $SS(L)$ is contained in $\Pi_{\mathbf{V},\mathbf{W}^{1}\oplus\mathbf{W}^{2}}$, it suffices to check for $L=L_{\boldsymbol{\nu}^{\prime}\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}}$. By applying the proof of [18, Theorem 13.3 and Corollary 13.6] to $Q^{(2)}$, we can see that any point in $SS(L_{\boldsymbol{\nu}^{\prime}\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}})$ admits a flag $f$ of flag type $\boldsymbol{\nu}^{\prime}\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}$. In the flag $f$, there are two special subspaces. The first one is $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{2}$ corresponding to $\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}$, the second one is $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}$ corresponding to $\boldsymbol{\omega}^{1}\boldsymbol{\nu}^{\prime\prime}\boldsymbol{\omega}^{2}$. In particular, $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{1}\oplus\mathbf{W}^{2}$ is $(x,y^{1},z^{1},y^{2},z^{2})$-stable, and so $(z^{1}+z^{2})(\mathbf{W})\subset\mathbf{V}^{\prime\prime}$. Similarly, $\mathbf{V}^{\prime\prime}\oplus\mathbf{W}^{2}$ is $(x,y^{1},z^{1},y^{2},z^{2})$-stable, and so $(y^{1}+y^{2})(\mathbf{V}^{\prime\prime})\subset\mathbf{W}^{2}$. Notice that $z^{2}=0$ implies that $(z^{1}+x^{2})(\mathbf{W}^{2})=0$. We get the proof. ∎
@@ 6.3. Characteristic cycles for framed quivers
By Proposition 6.4, $\mathcal{Q}_{\nu,\omega^{1}}$ is a full subcategory of $\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}})$, hence $\mathcal{K}(\omega^{1})_{-1}^{\psi}$ can be regarded as a sub $\mathbb{Z}$-module of $\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}}))$. Compose the characteristic cycle map with $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}]$, we obtain a morphism
@@
here $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}]$ is the linear map $\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}}))\rightarrow\mathcal{K}_{0}(\mathcal{D}^{b}_{\mathbf{G}_{\mathbf{V}}}(\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}},\Lambda_{\mathbf{V},\mathbf{W}^{1}}))$ induced by the functor $\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}:L\mapsto L[{\rm{dim}}\mathbb{T}]$.
@@
(7) $\mathbf{Ind}^{(r)}_{i}\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}\cong\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}\mathcal{F}^{(r)}_{i},$
and $[\textrm{For}_{\mathbf{G}_{\mathbf{V}}}^{\mathbf{G}_{\mathbf{V}}\times\mathbb{T}}]$ intertwines the operator $[\mathcal{F}^{(r)}_{i}]$ and the operator $[\mathbf{Ind}^{(r)}_{i}]$. Here we denote $\mathbf{Ind}$ by $\mathbf{Ind}^{(r)}_{i}$, in order to emphasize the change of dimension vectors.
Let $\bar{F}_{\mathbf{W}^{1}}$ be the subset of $T^{\ast}\mathbf{E}_{\mathbf{V}\oplus\mathbf{W}^{1},\Omega^{(1)}}$ consisting of $(x,y,z)$ such that $(x,y,z)(\mathbf{V}^{\prime\prime})\subseteq\mathbf{V}^{\prime\prime}$ and denote $\bar{F}_{\mathbf{W}^{1}}\cap\Lambda_{\mathbf{V},\mathbf{W}^{1}}$ by $\bar{F}^{nil}_{\mathbf{W}^{1}}$. By [14, Lemma 9.4], we have $\mathbf{G}_{\mathbf{V}}\times^{P}\bar{F}_{\mathbf{W}^{1}}\cong T^{\ast}_{W}(X\times X^{\prime})$, and we also have the following commutative diagram,
@@
hence the induction operator $\mathbf{Ind}^{(r)}_{i}:\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\Lambda_{\mathbf{V}^{\prime\prime},\mathbf{W}^{1}},\mathbb{Z})\rightarrow\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\Lambda_{\mathbf{V},\mathbf{W}^{1}},\mathbb{Z})$ is well-defined. By [14, Theorem 9.11], the top degree homology groups of CoHA considered in [31] and [9] realize the nilpotent enveloping algebra. Hence together with $[\mathbf{Ind}^{(r)}_{i}]^{\psi}$, the homology group $\bigoplus\limits_{\nu\in\mathbb{N}I}\mathbf{H}^{\mathrm{BM}}_{\mathrm{top}}(\Lambda_{\mathbf{V},\mathbf{W}^{1}},\mathbb{Z})$ becomes a ${}_{\mathbb{Z}}\mathbf{U}^{-}(\mathfrak{g})$-module. We have the following lemma.
@@
is a $\mathbf{U}^{-}(\mathfrak{g})$-linear map. Moreover, this map is independent on the choice of $\Omega$ with respect to Fourier-Sato transforms.
@@
By equation (7) and Theorem 2.2, the map ${\rm{CC}}^{\omega^{1}}$ intertwines the operator $[\mathcal{F}^{(r)}_{i}]$ and the operator $[\mathbf{Ind}^{(r)}_{i}]$. Since we use the same $\psi$-twist, the proposition is proved. By [17, Exercise IX.7], the characteristic cycle map commutes with Fourier-Sato transforms, hence ${\rm{CC}}^{\omega^{1}}$ is independent on the choice of $\Omega$. ∎
@@
by using functorial properties of characteristic cycles, where $\mathfrak{P}^{(r)}_{i}$ is the generalization of the Hecke correspondence in [27, Section 5.3]. Then ${\rm{CC}}^{s,\omega^{1}}$ is ${}_{\mathbb{Z}}\mathbf{U}(\mathfrak{g})$-linear directly, but we do not need such strong properties here.
Remark 6.19.
In [7], the authors consider the mixed Hodge modules on Grassmannians and construct an exact functor from the mixed Hodge modules to the $\mathcal{D}_{X,h}$-modules. The authors in [7] also point out that there should be a generalization to localized mixed Hodge modules on quivers, which could establish a relation between canonical bases and the quantized quiver varieties considered in [6] or the categorical realization via coherent sheaves in [8]. Indeed, the results in [7] solve this problem in $A_{1}$ quiver case.